OpenStudy (anonymous):

What number is 24% of 105? Please tell me how to solve w/o a calculator

OpenStudy (anonymous):

the answer is 25.2 but solving it with out a calculator Ummm

OpenStudy (agreene):

105*0.24 = 25.2 to do without calculator you just have to use long form multiplication and try not to mess up the decimal.

OpenStudy (anonymous):

Hmm, if I did 105x0.24 and round it would it still be an acceptable answer, like on a test or does that depend on the professor?

OpenStudy (anonymous):

I mean rounding to 25

OpenStudy (agreene):

depends on professor, in this case there is only one decimal so I would keep it. alternatively you can do fractions, if you are good with them. 0.24=6/25 105*6/25 = 126/5 which is an exact answer, also.

OpenStudy (anonymous):

Sorry, but how did you get 126?

OpenStudy (agreene):

\[\frac{105}{1}*\frac{6}{25} = \frac{105*6}{1*25}=\frac{630}{25}\] then take a factor of 5 out of both top and bottom (divide by 5) and you get \[\frac{126}{5}\]

OpenStudy (anonymous):

Oh thanks